I need help solving this problem I do not have a clue how to do it...thank you for you help in advance


Found 2 solutions by stanbon, malakumar_kos@yahoo.com:
Answer by stanbon(48558)

L has y-intercept (0, 2) and is perpendicular to the line with equation
2x - 3y =6.
---------
Find slope of the given line:
y = (2/3)x-2
slope of the given line is 2/3
-----------------
Therefore slope of every perpendicular line is (-3/2)
--------------
You now know the slope and the y-intercept of the line you want:
y = (-3/2)x + 2
==============
Cheers,
Stan H.

Answer by malakumar_kos@yahoo.com(315)


L has y-intercept (0, 2) and is perpendicular to the line with equation
2x - 3y =6.

Let the eq'n to the line L be y = Mx+c ........(1)
its y- intercept is (0,2)
it is perpendicular to 2x-3y = 6
to find the slope reduce it to std form -3y = -2x+6
y = 2x/3-2
slope of the given eq'n = 2/3
hence slope for the line L is M = -1/m = -1/2/3 = -3/2
to find the constant c we should substitute for y,M &x
y = Mx+c
2 = -3/2(0)+c or c = 2
Therefore the required eq'n is y = -3x/2+2
multiply by 2 throughout we get 2y = -3x+4